OpenStudy (anonymous):

if the equation x^3-hx^2+kx-9=0 has only one real root which has a value of 1, find the range of values of k.

OpenStudy (empty):

what can h be, just any real number?

OpenStudy (anonymous):

not mention

OpenStudy (empty):

Well here's how I solved it, just start with the factored form, since we know one of the roots is 1, but we also have two other roots too: \[(x-a)(x-b)(x-1)=x^3-hx^2+kx-9\] Now when we multiply it out on the left side we get: \[x^3-(a+b+1)x^2+(ab+a+b)x-ab=x^3-hx^2+kx-9\] now we can just set the coeffients equal to each other! \[h=a+b+1 \\ ab+a+b=k \\ ab=9\] Now look at the equation for k in the middle, we can subtract 1 from the first equation to get a+b nd we have ab from the second equation, so we can just plug them in like that \[9+(h-1)=k \]

OpenStudy (anonymous):

the answer is 3 < k <15

OpenStudy (rational):

\[x^3-hx^2+kx-9=0\] since \(1\) is a root, it must satisfy above eqn \[1-h+k-9 = 0\\h =k-8\] the eqn becomes \[x^3-(k-8)x^2 + kx -9\] The cubic will have only one root after and before the graph touches x axis. That is when the other root \(r\) has a multiplicity of \(2\) : \[x^3-(k-8)x^2 + kx -9 = (x-1)(x-r)^2\] Comparing constant terms we get \(r = \pm 3\) comparing \(x\) coefficients we get \(k = 2r+r^2\) use \(r\) values for getting the bounds of \(k\)